OpenStudy (poetryprincess):

If f(x) = x^3 – 2x^2, which expression is equivalent to f(i)? A.–2 + i B.–2 – i C.2 + i D.2 – i

OpenStudy (rational):

whats stopping you from plugging in x = i

OpenStudy (jhannybean):

Remember, \(i^2=-1\).

OpenStudy (jhannybean):

And also, \(i = \sqrt{-1}\)

OpenStudy (rational):

\[\large \begin{align}f(\color{red}{x}) &= \color{red}{x}^3-2\color{red}{x}^2\\~\\~\\f(\color{red}{i}) &=?\end{align}\]

OpenStudy (jhannybean):

\[(i)^3 = i^{2+1}=(i)^2 \cdot i\]

OpenStudy (poetryprincess):

Thanks a lot

OpenStudy (poetryprincess):

Both of you